An object of mass m = 0.500 kg is suspended from the ceiling of an accelerating truck. Taking a = 3.00 m/s 2 , find
(a) the angle u that the string makes with the vertical and
(b) the tension T in the string.
1
Expert's answer
2020-09-29T09:42:11-0400

"\tan u=F/mg=ma/mg=m/g\approx0.31"

"u=\arctan0.31\approx17^0"

"T=\sqrt{(mg)^2+(ma)^2}=m\sqrt{g^2+a^2}\approx5.1N"